OpenStudy (anonymous):

If a car engine burns an equivalent of 20,000 joules of energy and attains a kinetic energy of 5,000 J, how much energy is dissipated in the form of heat?

OpenStudy (anonymous):

I just don't know whether to add them or subtract them.

OpenStudy (anonymous):

\[\eta = \frac{ Eout }{ Ein } \times 100 % \] \[\eta = \frac{ 5000 }{ 20000} \times 100 percent\] \[\eta = 25 percent\]

OpenStudy (anonymous):

100 - 25 = 75 percent energy is dissipated in the form of heat :)

OpenStudy (anonymous):

Oh okay. So then it would be 15,000 J ? & Thank you for helping me!

OpenStudy (anonymous):

yes of course :) ur welcome

OpenStudy (anonymous):

dont forget about medal :p